OpenStudy (anonymous):

look at this pattern: 36, __, 16, 32/3, 64/9, 128/27 what is the missing number?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

starting from right to left we have \[\frac{2^7}{3^3},\frac{2^6}{3^2},\frac{2^5}{3^1}\]

OpenStudy (anonymous):

so going the other way looks like you are multiplying by 2 and dividing by 3

OpenStudy (anonymous):

\[36\times \frac{2}{3}=24,24\times \frac{2}{3}=16, 16\times \frac{2}{3}=\frac{32}{3}, ...\]

OpenStudy (anonymous):

so its 24?